solve log6(x) + log6(x-5)= 2
confused..

Answer by edjones(8007)

log6(x) + log6(x-5)= 2
6^log6(x) * 6^ log6(x-5)=6^2
x(x-5)=36
x^2-5x-36=0
(x-9)(x+4)=0
x=9
.
Ed